OpenStudy (anonymous):

4(x + 2) - 3(x - 5) ANSWER: x + 23

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

4x+8-3x+8

OpenStudy (callisto):

4(x + 2) - 3(x - 5) = 4(x) + 4(2) -3(x) -3(-5) = 4x + 8 - 3x +15 = 4x - 3x +8 + 15 = Can you continue?

OpenStudy (anonymous):

4(x+2) is 4x + 8, and -3(x - 5) is -3x + 5. so you have 4x + 8 - 3x + 15. 4x - 3x = x, 8 + 15 = 23. so x + 23, you're right, good